5. In a pediatric clinic a study was carried out to see how effective acetaminophen is in re- ducing temperature. Seven 5 year old children suffering from influenza had their tem- peratures taken immediately before and 1 hour after administration of acetaminophen. The results were: Patient Before After 102.4 99.6 103.2 100.1 101.9 100.2 103 101.1 101.2 99.8 100.7 100.2 102.5 101.0 (a) Use a test based on the normal assumption to determine whether the acetaminophen is effective in reducing temperature. i. State the hypotheses. ii. Evaluate the test statistic. iii. Calculate the P value. iv. Draw a conclusion in the context of the problem. V. Are the results statistically significant at the α .05 level? (b) Construct a 99% confidence interval for the difference in rneans
Answer #1

a) i) Two paired t-test:

patient Before After Diff
1 102.4 100 3
2 103.2 100.1 3
3 101.9 100.2 2
4 103 101.1 2
5 101.2 99.8 1
6 100.7 100.2 1
7 102.5 101 2
Mean 102.13 100.2857 1.842857
Std 0.9196 0.566947 0.879123

H_0:\mu _d=0 and H_a:\mu _d>0

Test statistic:

t= \frac{\bar{d}}{s/\sqrt{n}}= \frac{1.842857}{0.879123/\sqrt{7}}=5.546143

P-value: 0.001

The test statistic is significant and rejects H0. There is evidence to support the claim.

b) 99% confidence interval:

df= n-1= 6

\bar{d}\pm t_{df,\alpha /2}*\frac{s}{\sqrt{n}}=> 1.842857\pm 3.707*(\frac{0.879123}{\sqrt{6}})

Lower limit: 1.842857- 1.231752=0.611105

Upper limit: 1.842857+ 1.231752=3.074609